OpenStudy (anonymous):

Use the quadratic formula to find the zeros of the function. Round to tenths if necessary. y = 17x2 - 21 A. {-1.1, 1.1} B. {1.1} C. {-4.1, 4.6} D. no real solution

OpenStudy (hoblos):

A

OpenStudy (anonymous):

explain pls :)

OpenStudy (anonymous):

first equate the equation to 0

OpenStudy (anonymous):

substitute the values in the quadratic formula

OpenStudy (hoblos):

same formula given before with b=0

OpenStudy (anonymous):

\[\Large \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\] Knock yourselves out ;)

OpenStudy (anonymous):

that

OpenStudy (anonymous):

Remember a quadratic is a second-degree polynomial to the form ax^2 + bx + c = 0. Plug the appropriate values in the quadratic formula, presto!

OpenStudy (anonymous):

I always feel that formula is very lengthy to remember and use.

OpenStudy (hoblos):

you must get to it because you will always need it ;)

OpenStudy (anonymous):

I don't specifically remember when I first started memorising the quadratic formula, but what I can remember is the uncountable number of times it's been a useful addition to the memory bank. Like anything, use it a few dozen or so times and it's stuck there for life ;)

OpenStudy (anonymous):

(x+1.11) and (x-1.11) so let x+1.11=0 and x-1.11=0 then get x=-1.11 x=1.11

OpenStudy (anonymous):

Unfortunately, you guys are right, that formula's very tough but equally useful......